OpenStudy (anonymous):

Factor Completely: 2x^5-7x^3-4x

OpenStudy (anonymous):

I can get to (x-2) x (x+2) (2 x^2+1)

OpenStudy (anonymous):

see ur answer is following : 2x^5 - -8(x^3) + (x^3) - 4x 2x^3(x^2 - 4) + x(x^2 - 4) (2x^3 +x)(x^2-4) or we can say: (x+2)(x-2)x(2x^2+1) so the factors of the given expressions are: 1) (x+2) 2) (x-2) 3) (x) 4) (2x^2 + 1)

OpenStudy (anonymous):

x(2x^4 - 7x^2 - 4) =x(2x^2 + 1 )(x^2 -4) = x(x-2)(x+2)(2x^2 + 1)

OpenStudy (anonymous):

did u understand it kimbanaag

OpenStudy (anonymous):

yes thanks =)

OpenStudy (anonymous):

then give medal and become fan if U R SATISFIED

OpenStudy (anonymous):

thnx